OpenStudy (anonymous):

Factor completely: 3ab – 5ax + 9b − 15x

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

3ab-5ax+9b-15x = 3ab+9b-5ax-15x = 3b(a+3)-5x(a+3) = (3b-5x)(a+3)